Please help, this one is too much! Solve using the substitution method.
7x + y = 14 and 2x - 5y = -33

Answer by stanbon(75887)

solve the 1st equation for "y" to get:
y = 14-7x
Substitute that into the 2nd equation to get:
2x - 5(14-7x)= -33
2x - 70+35x = -33
37x = 37
x= 1
Substitute that into y=14-7x
to get y=14-7(1)
y = 7
So, the solution is x=1, y = 7
cheers,
Stan H.